# RuleArena: A Benchmark for Rule-Guided Reasoning with LLMs in Real-World Scenarios

### v1

@@ B.2 NBA Data Annotation
Our annotators are first required to create diverse valid scenarios involving one or more teams and players, as the following “team_situations” and “players_situations”, and provide the number of teams (“n_teams”) and players (“n_players”) involved. Each item in the “team_situations” list indicates the current salary of the team and its available first-round draft picks, while each item in the“player_situations” list tells the player’s information including his draft year, age, and current (or last) contract.
@@ Appendix C Structured Rule Extraction in Each Scenario
As introduced in Section 4.1, we utilize the structured output mode of GPT-4o (OpenAI, 2024) to convert LLMs’ textual output into structured data. Here we present the data structure we used in parsing.
@@ D.1 Rule-Wise Statistics
We visualize the rule-wise recall, precision, and correctness in Figure 6-8. Since precision is always $1.0$ in airline and tax domains, we skip these two charts.

### v2

@@ B.2 NBA Data Annotation
Our annotators are first required to create diverse valid scenarios involving one or more teams and players, as the following “team_situations” and “players_situations”, and provide the number of teams (“n_teams”) and players (“n_players”) involved. Each item in the “team_situations” list indicates the current salary of the team and its available first-round draft picks, while each item in the“player_situations” list tells the player’s information (i.e., draft year, age, and current (or last) contract). All players and teams are anonymized as Player (Team) A/B/C/… to avoid data leakage.
@@
To ensure the quality of annotation, we provide each annotator with a detailed annotation document and training sessions, and ask our annotators to annotate a small subset of problems and give explanations for verification before annotation. Only if each instance in the verification subset is correct, the annotator will be invited for the formal annotation.
@@ Appendix C Structured Rule Extraction in Each Scenario
As introduced in Section 4.1, we utilize the structured output mode of GPT-4o (OpenAI, 2024a) to convert LLMs’ textual output into structured data. Here we present the data structure we used in parsing.
@@ D.1 Rule-Wise Statistics
We visualize the rule-wise recall, precision, and correctness in Figure 4-6. Since precision is always $1.0$ in airline and tax domains, we skip these two charts.
@@
D.2 What Impacts Rule Following?
D.2.1 Correlation Between Accuracy and Other Metrics
(a) Airline
(b) NBA
(c) Tax
D.2.2 Do In-Context Examples Help?
Rule Setting ${\rm R}(r)$ ${\rm P}(r)$
Over 38 rule 0-shot 0.00 N/A
1-shot 0.35 0.76
[…]
1-shot 0.23 0.20
Table 8: 0-shot and 1-shot rule-wise comparison.
D.2.3 Does Rule Representation Matter?
Models Setting Airline Tax
${\rm AC}(t)$ ${\rm R}(t)$ ${\rm Acc}(t)$ ${\rm AC}(t)$ ${\rm R}(t)$ ${\rm Acc}(t)$
Llama 70B Table 0.764 0.558 0.01 0.834 0.989 0.01
Text 0.764 0.582 0.01 0.814 0.991 0.00
[…]
Text 0.748 0.633 0.02 0.859 0.996 0.01
[…]
Text 0.835 0.587 0.07 0.919 0.998 0.05
[…]
Text 0.937 0.705 0.06 0.971 1.000 0.33
[…]
Text 0.864 0.669 0.03 0.960 1.000 0.33
D.2.4 Do Distractive Rules Matter?
(a) Correctness
(b) Recall
(c) Accuracy
D.2.5 Can Tool Augmentation Help?
In Appendix D.2.1, we notice that only near-perfect application correctness ${\rm AC}(t)$ can lead to significant accuracy improvement. As the most simple way to reduce errors, especially in mathematical and logical operations, is to introduce external tools, we wonder to what extent external tools can help in our rule-guided reasoning tasks.
Following program of thoughts (Chen et al., 2023) prompting, we ask our LLMs to write Python code to calculate the answer on airline bag fee tasks by defining a solution() function and returning the total_cost variable, and the execution result of solution() function is viewed as the predicted answer. In this way, the Python interpreter can be viewed as an oracle tool for mathematical and logical calculation. To ensure the correct format of response, we use the 1-shot setting, so we compare the additional results with the original 1-shot results in Table 3 as follows:
Models Setting Level 1 Level 2
${\rm AC}(t)$ ${\rm R}(t)$ ${\rm Acc}(t)$ ${\rm AC}(t)$ ${\rm R}(t)$ ${\rm Acc}(t)$
Llama 70B 1-shot Default 0.809 0.787 0.17 0.827 0.801 0.07
Tool Augmented 0.863 0.882 0.34 0.827 0.887 0.18
[…]
Tool Augmented 0.939 0.899 0.42 0.946 0.899 0.26
[…]
Tool Augmented 0.939 0.914 0.44 0.937 0.940 0.33
Table 10: Results of different LLMs with tool augmentation on airline tasks.
As can be seen from these results, when provided with oracle math and logic tools, LLMs can achieve a significant performance boost in terms of accuracy ${\rm Acc}(t)$. However, even provided with such tools, LLMs are far from being able to resolve our rule-guided reasoning tasks. We observe non-perfect recall ${\rm R}(t)$ and correctness ${\rm AC}(t)$, which indicates that LLMs still make mistakes in generated codes.
D.2.6 Summary of Factors that Influence Rule-Guided Following
In summary, various factors, such as rule complexity, the presence of distractive information, and the difficulty gap between in-context examples and target problems, can profoundly influence LLM performance. Even when LLMs succeed in simpler conditions, challenges like complex mathematical reasoning, large amounts of extraneous rules, and non-ideal in-context samples can severely limit their effectiveness on RuleArena problems.